A) A solution contains 1.49×10-2 M potassium phosphate and 6.81×10-3 M potassium bromide.
Solid silver acetate is added slowly to this mixture.

What is the concentration of bromide ion when phosphate ion begins to precipitate?

B) A solution contains 1.22×10-2 M chromium(III) acetate and 1.22×10-2 M aluminum nitrate.
Solid sodium phosphate is added slowly to this mixture.

What is the concentration of chromium(III) ion when aluminum ion begins to precipitate?
Answer #1

A) Ksp of Ag3PO4 = [Ag+]3*[PO43-]

i.e. 8.89*10-17 = [Ag+]3*1.49*10-2

i.e. [Ag+]= 1.814*10-5 M

Therefore, the solubility of silver phsophate = 6.05*10-6 M

Hence, the concentration of bromide ion when phosphate ion begins to precipitate = 6.81*10-3 - 6.05*10-6 = 6.804*10-3 M

B) Ksp of AlPO4 = [Al3+]*[PO43-]

i.e. 6.3*10-19 = 1.22*10-2*[PO43-]

i.e. [PO43-]= 5.164*10-17 M

Therefore, the solubility of aluminium phsophate = 5.164*10-17 M

Hence, the concentration of chromium(III) ion when aluminium ion begins to precipitate = 1.22*10-2 - 5.164*10-17 = 1.22*10-2 M